Solve each system by graphing.
y = 2x + 1
X +Y = -2
: Solve each system by graphing.
y = 2x + 1
X +Y = -2

Answer by salvyb(3)

y=2x+1_X+Y=-2
Since Y does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting Y from both sides.
y=2x+1_X=-Y-2
This system is not in terms of x and y therefore it can't be solved using the graphing method.